The owner of a go cart track was counting the money from one days ticket sales
He knew that a total of 168 tickets were sold. Adult tickets cost 7.50 each and
childrens tickets cost 4.75 each. The total receipts for the day was 996.00 how many of each type of ticket was sold.

Answer by checkley71(8403)

A+C=168 OR A=168-C
7.50A+4.75C=996
7.50(168-C)+4.75C=996
1260-7.5C+4.75C=996
-2.75C=996-1260
-2.75C=-264
C=-264/-2.75
C=96 THE NUMBER OF CHILDREN'S TICKETS SOLD.
168-96=72 ADULT TICKETS SOLD.
PROOF:
7.50*72+4.75*96=996
540+456=996
996=996